OpenStudy (anonymous):

if y=Sin[2^x]. To find dy/dx using the chain rule (dy/dx = dy/du du/dx) we have u= what? A. Sin[x] B. 2^x C. (2^x)Ln[2]Cos[2^x]Sin[2^x] D. none of the above.

OpenStudy (anonymous):

you are trying to solve this using substitution and they are asking you what will you use for u?

OpenStudy (anonymous):

is it 2^x?

OpenStudy (anonymous):

reo.j.syhat: that's correct

OpenStudy (anonymous):

reo is right that will give you sin(u)

OpenStudy (anonymous):

ok, on the inside function ur going to use d/dx(a^x) = lna(a^x)

OpenStudy (anonymous):

for the outside function, recognize that d/dx(sinx) = cosx

OpenStudy (anonymous):

so c would be correct, i think